OpenStudy (spicytaco1):

Multiply. Express your answer in lowest terms. 4/5 x 3/8 =

OpenStudy (anonymous):

\[\frac{ 4 }{ 5 }*\frac{ 3 }{ 8 }=\frac{ 12 }{ 40 }\]

OpenStudy (anonymous):

Then simplify 12/40